OpenStudy (anonymous):

Determine the potential difference between two charged parallel plates that are 0.20 cm apart and have an electric field strength of 6.0 V/cm.

OpenStudy (anonymous):

E = V / s s is separation 6.0 V/cm = V /0.2 cm solve for V